ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57, 58 CHO, Yong-Seung, Ewha Women’s University, (South)Korea Finite Group Actions and Seiberg-Witten Invariants. Let L be a complex line bundle over a closed, oriented, smooth four-manifold X such that the first Chern class of L is equal to the second Stiefel-Whitney class of the tangent bundle TX modulo two. Let a finite group G act on X as orientation preserving isometries and on L such that the projection from L to X is a G-map. We investigate the G-action on the Seiberg-Witten equations for the bundle L and the relations between the G-invariant Seiberg-Witten invariants on X and the ones on its quotient setting. References: Y. S. Cho, Finite Group Actions on SpinC Bundles, Preprint. Y. S. Cho, Finite group actions on the moduli space of self-dual connections I, Transactions of A. M. S. 323 (1991) 233-261.